QUESTION 5

  1. The marginal product for labor is given (MP) = 3 – 0.02*L; price of the product is $100 and wage = 200.  Based on information above, the marginal product of labor at the optimal level of employment is

    $3

    $2

    $1.5

    $1

2 points   

QUESTION 6

  1. If the labor elasticity of output is 0.5 and the capital elasticity of output is 0.9, then the production function exhibits

    constant returns to scale.

    economies of scale.

    diseconomies of scale.

    diminishing returns.

2 points   

QUESTION 7

  1. Suppose the long run production function is given by: Q = 4*L +2K2. Marginal product of labor (MPL) = 4 and wage is $10. Marginal product of capital (MPK) = 4K and price of capital (K) is $10. Consider the allocation labor (L) = 10 and capital (K) = 2. Based on information, the MRTS is equal to

    4

    2.5

    1

    0.5

2 points   

QUESTION 8

  1. If the demand for product increases,

    labor demand increases.

    labor demand decreases.

    labor supply decreases.

    labor supply increases.

2 points   

QUESTION 9

  1. Suppose a firm is operating in both a perfectly competitive product market and perfectly labor market. The firm’s short run production is Q = L2; where Q is output and L is labor, expressed in millions. Marginal product of labor (MPL) = 2L and wage is 10. The price of the product is $ 2. Based on information, the short run optimal level of employment is

    4 million

    2.5 million

    5 million

    0.4 million

2 points   

QUESTION 10

  1. Consider the following production function: Q = KL where Q = output, L = labor and K = capital. The marginal product of labor is given by MPL = K while the marginal product of capital is given by MPK = L.   If L = 10 and K= 5, the marginal product of capital is

    2

    5

    10

    50
Answer #1

Ans 5 )

OPTION- B

At optimal level of output,

Marginal revenue product of labor (MRPL) = Wage

MRPL = Price × marginal labor => 100 × (3 - 0.02L)

MRPL = 300 - 2L

200 = 300 - 2L

2L = 100

L = 100/2 = 50

Marginal product of labor = 3 - 0.02(50) = 3 - 1 = $2

Ans 6.) Option- B

Labor Elasticity of Output (EL) = 0.5

Capital elasticity of output (EC) = 0.9

EL + EC = 0.5 + 0.9 = 1.4

As the sum of both the elasticities is greater than 1, so, the production function exhibits increasing returns ro scale.

Increasing returns means that the cost of production decreases as more output is produced because each additional unit of input is able to produce more output.

Ans 7.) OPTION - D

MRTS = Marginal productLabor / Marginal productCapital

MRTS = 4/4K = 1/K

As, K = 2

MRTS = 1/2 = 0.5

  

Ans 8.) Option - B

If the demand for a good decreases,the demand for labor also decreases as firms need less labor to produce the good.